# Statistical Indications of Toponium Formation in Top Quark Pair Production

### v1

@@ Abstract
We present an analysis of six differential cross-section measurements of top-quark pair production in the dilepton channel from the ATLAS and CMS experiments. The data are compared to state-of-the-art QCD predictions with and without the inclusion of toponium formation effects. This contribution is modelled via a re-weighting of fixed-order matrix elements using the Green’s function of the non-relativistic QCD Hamiltonian, and we employ a statistical model to quantify the preference of the data for the toponium hypothesis. All observables yield Bayes factors larger than unity, with two exceeding 20, yielding strong evidence for toponium formation in top-quark pair production at the LHC.
@@
In the SM, top-quark pairs produced near threshold in a colour-singlet configuration can experience Coulombic gluon exchange, leading to an enhancement of the production cross section relative to the conventional, non-resonant process. This phenomenon can be interpreted as toponium formation whose effects are calculable using the Green’s function of the non-relativistic QCD Hamiltonian Fadin:1987wz ; Fadin:1990wx ; Hagiwara:2008df ; Sumino:2010bv . In particular, toponium formation is expected to modify both the invariant mass distribution and the angular correlations of the $t\bar{t}$ system and its decay products Fuks:2021xje ; Fuks:2024yjj . In this context, precision fixed-order computations have been achieved for specific observables Ju:2020otc ; Garzelli:2024uhe ; Nason:2025hix , while recent development of Monte Carlo simulations including toponium effects complementarily enable predictions for a wide range of particle-level observables at the LHC Fuks:2021xje ; Fuks:2024yjj ; Fuks:2025wtq .
@@
The SM predictions for conventional $t\bar{t}$ production and decay, referred to as the null hypothesis $\text{H}_{0}$ in our analysis, are obtained from fixed-order NNLO calculations as provided in Czakon:2020qbd , using the NNLO set of NNPDF3.1 parton distribution functions Bertone:2017bme . We then define our alternative hypothesis $\text{H}_{\mathrm{toponium}}$ as the coherent sum of the conventional and toponium contributions. The impact of the latter is simulated following the prescription of Fuks:2024yjj , implemented in MG5_aMC v3_5_7 Alwall:2014hca and in which the matrix element is re-weighted by the Green’s function of the non-relativistic QCD Hamiltonian. Parton-level events are then passed to Pythia v8.312 Bierlich:2022pfr for parton showering and hadronisation. Next, event selection and observable definitions follow those of the ATLAS and CMS measurements as implemented within the Rivet v4.1.0 framework Bierlich:2024vqo , using the corresponding public routines identified by the Rivet codes ATLAS_2019_I1759875 and CMS_2018_I1703993.
@@
where $\mathrm{d}X$ denotes the vector of absolute differences between the data and the prediction and $V$ is the data covariance matrix. The covariance matrices for the CMS results are taken from the Hepdata database hepdata.89307.v1 , while as covariance matrices for the ATLAS results are not publicly available, a diagonal covariance matrix is applied instead. A likelihood-ratio test statistic is constructed as $\lambda=e^{-\frac{1}{2}\Delta\chi^{2}}$, with $\Delta\chi^{2}$ being the difference between the $\chi^{2}$ values computed for the conventional SM hypothesis $\text{H}_{0}$ and the alternative hypothesis $\text{H}_{\mathrm{toponium}}$ including toponium contributions. The distributions of $\lambda$ under $\text{H}_{0}$ and $\text{H}_{\mathrm{toponium}}$ are then obtained via pseudo-experiments assuming Gaussian data and using the covariance matrix to model inter-bin correlations. The code used for the analysis is publicly available. [^1] The codebase is released under a GNU General Public License at GitHub.
@@
The extracted $\hat{\sigma}_{\rm toponium}$ values are broadly consistent across observables within one standard deviation, though they tend to exceed both the theoretical prediction and the ATLAS measurement. This pattern may reflect missing higher-order corrections in the conventional predictions, NNLO effects in top-quark decay not included in the LHC analyses, differences in fiducial phase-space definitions or new physics contributions as explored in Flacke:prep . Future particle-level measurements optimised for toponium sensitivity, released with public Rivet routines and associated covariance matrices as well as higher-order QCD and electroweak predictions, will allow these mild tensions to be probed with greater precision.
@@
Summary and Conclusions – We have analysed six particle-level differential cross sections of $t\bar{t}$ production measured by the ATLAS and CMS collaborations to probe for evidence of toponium formation. The data were compared to conventional predictions at NNLO QCD accuracy and to predictions incorporating toponium contributions. In all observables, the inclusion of toponium formation improves the agreement between data and theory. In addition, a set of one-parameter $\chi^{2}$ minimisations was performed to extract the inclusive toponium production cross section associated with each distribution. The resulting estimates are broadly consistent across observables, though they show mild tension with both the theoretical prediction and a recent detector-level measurement from ATLAS. Taken together, the improved modelling and consistent cross-section estimates provide evidence for toponium formation in LHC data. However, a combined determination of the significance and inclusive cross section across all observables is not currently feasible owing to the absence of publicly available information on correlations between distributions and between ATLAS and CMS data. A joint treatment of the two datasets, or a public release of inter-distribution and inter-experiment covariance matrices, would enable a definitive combined result and provide deeper insight into the presence of toponium in data.

### v2

@@ Abstract
We present an analysis of six differential cross-section measurements of top-quark pair production in the dilepton channel from the ATLAS and CMS experiments. The data are compared to state-of-the-art QCD predictions with and without the inclusion of toponium formation effects. This contribution is modelled via a re-weighting of fixed-order matrix elements using the Green’s function of the non-relativistic QCD Hamiltonian, and we employ a statistical model to quantify the preference of the data for the toponium hypothesis. All observables yield Bayes factors larger than unity, with two exceeding 20, yielding strong evidence for the toponium hypothesis in top-quark pair production at the LHC.
@@
In the SM, top-quark pairs produced near threshold in a colour-singlet configuration can experience Coulombic gluon exchange, leading to an enhancement of the production cross section relative to the conventional, non-resonant process. This phenomenon can be interpreted as toponium formation whose effects are calculable using the Green’s function of the non-relativistic QCD Hamiltonian Fadin:1987wz ; Fadin:1990wx ; Hagiwara:2008df ; Sumino:2010bv . In particular, toponium formation is expected to modify both the invariant mass distribution and the angular correlations of the $t\bar{t}$ system and its decay products Fuks:2021xje ; Fuks:2024yjj . In this context, precision fixed-order computations have been achieved for specific observables Ju:2020otc ; Garzelli:2024uhe ; Nason:2025hix , while recent development of Monte Carlo simulations including toponium effects complementarily enables predictions for a wide range of particle-level observables at the LHC Fuks:2021xje ; Fuks:2024yjj ; Fuks:2025wtq .
@@
The SM predictions for conventional $t\bar{t}$ production and decay, referred to as the null hypothesis $\text{H}_{0}$ in our analysis, are obtained from fixed-order NNLO calculations as provided in Czakon:2020qbd , using the NNLO set of NNPDF3.1 parton distribution functions Bertone:2017bme . We then define our alternative hypothesis $\text{H}_{\mathrm{toponium}}$ as the coherent sum of the conventional and toponium contributions. The impact of the latter is simulated following the prescription of Fuks:2024yjj , implemented in MG5_aMC v3_5_7 Alwall:2014hca , in which the leading-order matrix element for the $gg\to b\bar{\ell}\nu_{\ell}\bar{b}\ell^{\prime}\bar{\nu}_{\ell^{\prime}}$ process is re-weighted by the Green’s function of the non-relativistic QCD Hamiltonian computed from a tree-level Coulomb potential. Parton-level events are then passed to Pythia v8.312 Bierlich:2022pfr for parton showering and hadronisation. Next, event selection and observable definitions follow those of the ATLAS and CMS measurements as implemented within the Rivet v4.1.0 framework Bierlich:2024vqo , using the corresponding public routines identified by the Rivet codes ATLAS_2019_I1759875 and CMS_2018_I1703993.
@@
where $\mathrm{d}X$ denotes the vector of absolute differences between the data and the prediction and $V$ is the data covariance matrix. The covariance matrices for the CMS results are taken from the HepData database hepdata.89307.v1 ; for the ATLAS case, we instead adopt diagonal covariances matrices based on the bin uncertainties distributions. Although correlation matrices are provided in hepdata.91242 , we refrain from converting these into covariance matrices to avoid potential errors associated with subtleties in the definition of the correlation matrices. We subsequently encourage direct publication of covariance matrices for differential cross section data. A likelihood-ratio test statistic is constructed as $\lambda=e^{-\frac{1}{2}\Delta\chi^{2}}$, with $\Delta\chi^{2}$ being the difference between the $\chi^{2}$ values computed for the conventional SM hypothesis $\text{H}_{0}$ and the alternative hypothesis $\text{H}_{\mathrm{toponium}}$ including toponium contributions. The distributions of $\lambda$ under $\text{H}_{0}$ and $\text{H}_{\mathrm{toponium}}$ are then obtained via pseudo-experiments assuming Gaussian data and using the covariance matrix to model inter-bin correlations. The code used for the analysis is publicly available. [^1] The codebase is released under a GNU General Public License at GitHub.
@@
The extracted $\hat{\sigma}_{\rm toponium}$ values are broadly consistent across observables within one standard deviation, though they tend to exceed both the theoretical prediction and the ATLAS measurement. This pattern may reflect missing higher-order corrections in the conventional predictions, NNLO effects in top-quark decay not included in the LHC analyses or subtle differences in fiducial phase-space definitions. The potential for new physics explanations, as explored for instance in Banik:2023vxa ; Coloretti:2023yyq ; Maltoni:2024wyh ; Lu:2024twj ; Llanes-Estrada:2024phk ; Djouadi:2024lyv , should therefore only be assessed after accounting for both these effects and toponium contributions, following the path described in Flacke:2025dwk . Future particle-level measurements optimised for toponium sensitivity, released with public Rivet routines and associated covariance matrices as well as higher-order QCD and electroweak predictions, will allow these mild tensions to be probed with greater precision.
@@
Summary and Conclusions – We have analysed six particle-level differential cross sections of $t\bar{t}$ production measured by the ATLAS and CMS collaborations to probe for evidence of toponium formation. The data were compared to conventional predictions at NNLO QCD accuracy and to predictions incorporating toponium contributions. In all observables, the inclusion of toponium formation improves the agreement between data and theory. In addition, a set of one-parameter $\chi^{2}$ minimisations was performed to extract the inclusive toponium production cross section associated with each distribution. The resulting estimates are broadly consistent across observables, though they show mild tension with both the theoretical prediction and a recent detector-level measurement from ATLAS. Taken together, the improved modelling and the consistency of the extracted cross-section estimates suggest that non-relativistic threshold effects beyond conventional perturbative predictions play an important role in describing LHC data. However, a combined determination of the significance and inclusive cross section across all observables is not currently feasible owing to the absence of publicly available information on correlations between distributions and between ATLAS and CMS data. A joint treatment of the two datasets, or a public release of inter-distribution and inter-experiment covariance matrices, would enable a definitive combined result and provide deeper insight into the presence of toponium in data.